### v1

# Seeing Clearly, Answering Incorrectly: A Multimodal Robustness Benchmark for Evaluating MLLMs on Leading Questions

Abstract
Multimodal Large Language Models (MLLMs) have exhibited impressive capabilities in visual understanding and reasoning, providing sightly reasonable answers, such as image descriptions. This has spurred extensive research on the evaluation of MLLMs. Most evaluation benchmarks assume that incorrect answers indicate a lack of understanding of the visual content. However, our findings reveal that, in many cases, MLLMs answer questions incorrectly despite correctly understanding the visual content. This suggests that incorrect answers do not necessarily imply a lack of comprehension but may instead result from lacking robustness to leading questions. To comprehensively measure MLLMs’ understanding capability and robustness to leading questions, we introduce a MultiModal Robustness benchmark (MMR). MMR contains paired positive and negative questions across 12 categories, meticulously annotated by humans. We evaluate 18 leading MLLMs on the MMB benchmark, revealing that MLLMs suffer from fragility to leading questions despite understanding the visual content. To enhance MLLMs’ understanding capability and robustness, we further present a training set with paired positive and negative visual question-answer samples. Experiments verify that MLLMs’ robustness can be significantly enhanced by tuning on this new training set. The benchmark, training set, and code can be found at https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark.
1 Introduction
Multimodal Large Language Models (MLLMs) have garnered significant attention for their impressive capabilities in tackling diverse multimodal tasks, making notable strides in visual comprehension and reasoning [1; 2; 3]. This has spurred the development of various evaluation methods to assess the capabilities of MLLMs across different dimensions. These evaluation methods primarily encompass general benchmarks and hallucination benchmarks. General evaluation frameworks, such as MMBench [4], MMMU [5], VQA-v2 [6], GQA [7], ScienceQA [8], and et al., scrutinize models for their proficiency in generic and knowledge-based question and answer (VQA) tasks. In contrast, hallucination evaluation methods focus on evaluating the discrepancy between generated responses and the visual content. These methods typically add interference to the text (e.g., object [9; 10; 11; 12; 13; 14; 15; 16; 17], knowledge [9; 14; 15; 17], relation [18; 19; 15; 16], attribute [18; 20; 13; 14; 15; 16; 17] interference) to see if the MLLM can generate responses that are consistent with the visual content.
Figure 1: Left: Examples show that MLLMs understand content but answer incorrectly. Right: Performance comparison on positive and negative questions for the same visual content.
Most MLLM evaluation methods assume that incorrect answers indicate a lack of understanding of the visual content. However, we unveil a critical and overlooked phenomenon: MLLMs may produce incorrect answers despite correctly comprehending the visual content (as shown in Fig. 1). This suggests that MLLMs are susceptible to subtle interference on the question, hindering their ability to leverage their knowledge effectively. However, existing benchmarks and evaluation metrics, primarily focused on answer accuracy, fail to assess this phenomenon adequately. In this paper, we aim to construct a new benchmark for 1) assessing whether the model understands the visual content and 2) evaluating the robustness of MLLMs to leading questions.
To this end, we introduce the MultiModal Robustness benchmark (MMR). Each sample is manually labeled with one of 12 types of questions, with paired positive and leading questions. We design multiple-choice questions with answers and intrusive explanations to prevent the model from guessing. To quantify the evaluation, we devise a novel metric to measure the likelihood of a model’s failure when induced by a leading question, despite its comprehension of the visual content.
We evaluate 18 leading MLLMs on the MMR benchmark, as shown in Tab. 2. Surprisingly, these MLLMs perform unsatisfactorily in terms of leading questions. We attribute this issue to the scarcity of robustness training data. Previous datasets primarily consist of positive questions, with only a few including negative samples (such as LRV [9]) aiming at mitigating object hallucination. However, most existing training data is insufficient to effectively enhance the robustness of MLLMs, as shown in Tab. 3. To address this challenge, we introduce a data construction pipeline. We first extract visual information from images, including text/numbers, objects, attributes, relationships, and local/global comprehension. Based on this information, we generate paired positive and negative samples.
In summary, the contributions are three folds: 1) We introduce a human-annotated multimodal robustness benchmark (MMR) and new evaluation metrics for measuring MLLMs’ understanding capability and robustness to leading questions. 2) We present a data construction pipeline to enhance the understanding ability and robustness by generating paired positive and negative samples. 3) We extensively experiment on the MMR and general benchmarks. Results show that current MLLMs are severely affected by leading questions, which our proposed dataset can effectively mitigate.
2 Related Works
Multimodal large language models (MLLMs).
Building on the success of Large Language Models, recent research has increasingly focused on MLLMs [21; 1]. This shift aims to achieve better cross-modality understanding and reasoning. Currently, a significant amount of research is exploring various aspects of MLLMs, including the design of structures [22; 23], data construction [24; 25; 26], training strategies [27; 28; 29], and lightweight MLLMs [3; 30]. For a comprehensive review of these topics, readers are encouraged to investigate surveys [31; 32; 33] dedicated to several areas.
Evaluations of MLLMs.
The assessment of the MLLM’s capabilities is mainly divided into two main categories: general capability [8] and hallucination benchmarks [34]. General capability benchmarks can be divided into single-task benchmarks and comprehensive benchmarks. Single-task benchmarks, such as VQA [6], OK-VQA [35], VQA-v2 [6], GQA [7], and MS-COCO [36], are designed to assess the performance of models on a specific defined task. Comprehensive benchmarks, such as MMBench [4], SEED [37], MMMU [5], and MMStar [38] aim to assess MLLMs’ general multi-modal perception (such as color, position, and celebrity recognition.) and reasoning capabilities (such as numerical calculation and knowledge-based reasoning). Hallucination benchmarks mainly focus on evaluating the discrepancy between text responses generated by MLLMs and provided visual content [34]. Most hallucination benchmarks focus on object [39; 9; 10; 11; 12; 13; 14; 15; 16; 17], knowledge [9; 14; 15; 17], relation [18; 19; 15; 16], attribute [18; 20; 13; 14; 15; 16; 17], spurious images [40; 41] hallucinations. However, current benchmarks assume correct answers indicate understanding, while incorrect answers do not. We find that MLLMs answer incorrectly despite understanding the visual content.
Multimodal instruction-tuning data generation.
The generation of synthetic data for multimodal instruction tuning can be divided into two main methods: text-only-based and image-based generation. Text-only methods include LLaVA [22], which uses the text capabilities of GPT-4 [42] to build a multimodal dataset, InstructBLIP [43], which integrates VQA datasets focused on academic tasks to enhance visual comprehension, and LRV [9], which generates positive and negative samples using image bounding boxes for robust tuning. Image-based methods employ multimodal large language models to generate caption data [44; 24], VQA [7; 8; 45] complex reasoning instruction data [46; 47; 3; 48] and counterfactual data to address specific scenarios [18]. SPHINX-X [49] assembles an OCR-focused dataset from large-scale PDFs, converting text annotations and bounding boxes into a unified question-answering format.
3 MMR Benchmark
Benchmark Design.
As shown in Fig. 2, we design paired positive and negative questions for each sample. Each question has four brief options, with only one correct answer. The positive questions assess the model’s ability to understand the visual content correctly, while the negative questions evaluate the model’s robustness in answering correctly despite misleading prompts.
To achieve a comprehensive evaluation, we manually construct 300 positive and 300 leading negative questions across three levels: character, attribute, and context. Character-level questions prompt identifying elements like characters or numbers, while attribute-level questions focus on properties such as color, texture, and quantity. Context-level inquiries delve into higher-level concepts like emotions, culture, and common sense. The positive questions aim to evaluate the model’s understanding ability, while the misleading ones challenge its resistance to interference. For instance, character-level misleading questions alter elements like characters or numbers, attribute-level ones confuse properties, and context-level prompts introduce complex concepts, thoroughly testing the model’s resilience against misleading information.
Figure 2: Examples in the MMR benchmark. The colors    ,    ,    , denote character, attribute, and context-level, respectively. POS and NEG denote the positive and negative questions, ANS
denotes the answer.
Evaluation metrics.
We categorize the results into four groups: 1) Understanding and Robustness (UR), 2) Understanding but Fragility (UF), 3) not Understanding and Rigorous (NR), and 4) not Understanding and Fragility (NF). To quantify the effect of negative questions in misleading the model despite its comprehension of visual content, we introduce the “Misleading Rate (MR)”. Additionally, we introduce the “Robustness Accuracy (RA)” to evaluate the actual understanding capability of MLLMs. The formulation is as follows:
$\text{MR}=\frac{N_{UF}}{N_{UR}+N_{UF}}$ (1)
$\text{RA}=\frac{N_{UR}}{N_{UR}+N_{UF}+N_{NR}+N_{NF}}$ (2)
Where $N_{\text{i}}$ ($\text{i}\in\{UR,UF,NR,NF\}$) represents the number of samples. By combining these two evaluation metrics, our benchmark can not only reflect the model’s basic understanding capabilities but also reveal its robustness to misleading prompts.
4 MMR-data Construction
Table 1: Comparison between different instruction tuning datasets in MLLMs.
Characteristics LLaVA ALLaVA LRV Ours
Hard negative instructions?
Paired positive-negative data?
Information extraction?
Average conversation rounds 2.35 1 1 6.82
Figure 3: Overview of data generation. 1) Key information within the image is extracted using GPT-4V. 2) The extracted key information is then used to generate paired positive and negative samples. 3) Filtering low-quality samples.
To enhance MLLMs’ understanding capability and robustness, we propose a data construction pipeline that utilizes GPT-4V to generate paired positive and negative samples for instruction tuning, as shown in Fig. 3. The distinctions between the proposed dataset and previous conversation datasets are illustrated in Tab. 1.
Information extraction. Previous approaches that generate instruction-tuning data for multimodal conversations with GPT-4V primarily fall into two categories: direct generation and annotation-driven generation. Direct generation methods, such as ALLaVA [46], rely entirely on the basic ability of GPT-4V to synthesize samples. In contrast, annotation-driven generation methods, such as LRV [9], utilize existing image annotations (e.g.object bounding boxes and textual descriptions) as additional information for text-only GPT-4 to generate conversational samples. However, these methods fail to fully utilize the fine-grained details. In contrast, our data construction pipeline adopts a different approach by implicitly and comprehensively extracting this detailed information directly from the image. This includes 1) text or numbers (if present), 2) objects and people, 3) object attributes (e.g., colors, textures, locations) and human characteristics (e.g., postures, dresses), 4) interrelationships between people and objects, 5) events or activities, and 6) the overall feeling and understanding evoked by the image. The detailed prompts are in the Appendix.
Instruction tuning data generation. The most relevant work to our data generation approach is LRV [9], which uses the Visual Genome [50] dataset to guide GPT-4 in generating positive and negative samples about object presence and manipulation knowledge. However, LRV has two significant limitations: limited fine-grained information and unpaired negative samples. It relies solely on bounding boxes and object labels, omitting textual descriptions, numerical attributes, and deeper image comprehension. Additionally, LRV’s negative samples are not explicitly paired with positive counterparts, reducing training effectiveness. To address these challenges, we generate positive samples using extracted information and construct negative samples that directly contradict the positive ones. This ensures a strong contrast for more effective model training, providing richer context and paired positive and negative samples. Detailed prompts are in the Appendix.
Sample filtering. We collect part of COCO [51] images from the LLaVA 158k and construct paired positive and negative samples. We filter the samples by keyword matching, removing conversations with uncertain answers (e.g., “uncertain”) and redundant phrases (e.g., “in the image”).
5 Visual Instruction Tuning
Framework.
The structure consists of a visual encoder, a visual-language connector, and a language model. Specifically, we employ SigLIP [52] as the visual encoder, Phi-2 (2.7B) [53] as the language model, and a two-layer MLP as the connector. We build the model based on Bunny [3].
Training.
Given an image and text, the image is processed through a visual encoder to obtain visual features. These features are then adjusted via a connector to align with the dimensions of the language model. The text is tokenized to generate textual features. These visual and textual features are concatenated and fed into the language model for generating responses. During training, each sample comprises instruction and response. The instructions are masked, and only the response and the model’s output are used to calculate the loss.
We employ a two-stage training strategy. In the first stage, the visual and language alignment stage, we use Bunny-pretrain-LAION-2M [3] data to train only the connector while freezing the vision encoder and the language model. In the second stage, we fine-tune both the connector and the language model using the generated MMR dataset (MMR-data). To further enrich the dataset’s diversity, we include OCR data (OCR-VQA [54]), visual question-answering data (VQA-v2 [6], GQA [7], OK-VQA [55], A-OKVQA [35], Visual Genome [50], and RefCOCO [56]). Additionally, to maintain the language capabilities of the MLLM, we incorporated text-only data from WizardLM [57].
Table 2: Comparison to state-of-the-art MLLMs on the MMR benchmark. §We evaluate the officially released checkpoint by ourselves. Abbreviations: Char/Num. (Character/Number), Pres. (Presence), Color/Tex. (Color/Texture), Num. (Number), Shape (Shape), Posture (Posture), Pos. (Position), Abstract. (Abstract Knowledge), Concrete. (Concrete Knowledge), Expert (Expertise), Act. (Activity), Rel. (Relationships). $\uparrow$: higher is better, $\downarrow$: lower is better.
Method Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. RA $\uparrow$
Random 25 25 25 25 25 25 25 25 25 25 25 25 25
Closed-source Models
GPT4o [42] 72.50 68.18 66.67 45.83 87.5 70.83 50.00 68.18 76.19 70.97 83.33 63.64 69.00
Gemini-pro-vision [1] 42.50 50.00 41.67 25.00 83.33 50.00 45.45 40.91 47.62 45.16 70.83 45.45 48.67
Qwen-VL-max [58] 67.50 72.73 66.67 41.67 79.17 62.5 63.64 77.27 80.95 61.29 79.17 72.73 68.33
Claude3-Opus-V [59] 35.00 22.73 12.50 16.67 33.33 16.67 22.73 45.45 33.33 25.81 37.50 40.91 28.67
Step-1V [60] 60.00 54.55 58.33 20.83 70.83 54.17 31.82 54.55 57.14 45.16 79.17 50.00 53.33
Glm-4V [61] 60.00 54.55 54.17 29.17 58.33 41.67 27.27 72.73 47.62 35.48 70.83 45.45 50.00
Open-source Models
LLaVA-1.6-34B [62] 75.00 68.18 66.67 41.67 79.17 54.17 72.72 81.81 71.42 64.52 79.17 68.18 68.67
Mini-Gemini-HD-34B [23] 62.50 63.64 70.83 54.17 79.17 62.50 72.73 86.36 85.71 54.84 19.17 68.18 69.00
Yi-VL-34B [63] 52.50 63.64 70.83 41.67 75.00 37.50 59.09 68.18 57.14 48.39 70.83 63.64 58.33
InternVL-Chat-V1-5 (26B) [64] 62.5 59.09 66.67 41.67 66.67 41.67 54.55 63.64 66.67 45.16 79.17 72.73 59.67
MiniCPM-Llama3-V [65] 37.5 45.45 50.00 16.67 41.67 37.5 36.36 68.18 33.33 29.03 41.67 54.55 40.33
Cogvlm2-llama3 [66] 60.00 63.64 54.17 37.5 70.83 33.33 40.91 50.00 85.71 41.94 62.50 50.00 54.00
Idefics2-8B [67] 57.50 59.09 54.17 50.00 79.17 41.67 27.27 77.27 76.19 45.16 75.00 40.91 56.67
Bunny-Llama-3-8B-V [3] 55.00 63.64 54.17 37.50 79.17 62.50 54.55 72.73 85.71 48.39 75.00 50.00 60.67
LLaVA-RLHF (7B) [68] 7.50 36.36 33.33 33.33 50.00 16.67 9.09 59.09 38.10 22.58 50.00 31.82 30.67
Deepseek-VL-7B-Chat [29] 52.50 54.55 54.17 37.5 62.5 25.00 18.18 54.55 52.38 35.48 75.00 50.00 47.67
Mplug-owl2-llama2-7B [69] 32.50 63.64 58.33 20.83 62.50 37.50 13.64 54.55 47.62 25.81 58.33 31.82 41.33
Phi-3-vision (4B) [70] 62.50 59.09 58.33 37.50 70.83 33.33 31.82 54.55 66.67 41.94 58.33 50.00 52.33
Bunny-MMR-3B 60.0 59.09 58.33 25.0 83.33 50.0 54.55 68.18 57.14 51.61 79.17 54.55 58.33
Method Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
Random 50 50 50 50 50 50 50 50 50 50 50 50 50
Closed-source Models
GPT4o [42] 9.38 16.67 23.81 26.67 4.55 19.05 38.89 28.57 15.79 24.14 13.04 22.22 19.46
Gemini-pro-vision [1] 29.17 31.25 41.18 45.45 13.04 40.00 33.33 52.63 44.44 48.15 19.05 23.08 34.82
Qwen-VL-max [58] 22.86 11.11 23.81 28.57 5.00 25.00 30.00 19.05 19.05 29.63 9.52 15.79 20.23
Claude3-Opus-V [59] 26.32 50.00 72.73 55.56 50.00 71.43 50.00 37.5 46.15 60.00 43.75 30.77 48.50
Step-1V [60] 14.29 25.00 26.32 61.54 5.56 40.91 61.11 33.33 33.33 44.00 9.52 21.43 30.43
Glm-4V [61] 27.27 36.84 35.00 56.25 33.33 52.38 53.85 20.00 47.37 57.69 19.05 37.50 38.78
Open-source Models
LLaVA-1.6-34B [62] 6.25 11.76 20.00 23.08 9.52 35.00 11.11 14.28 25.00 20.00 9.52 16.67 16.26
Mini-Gemini-HD-34B [23] 21.88 12.50 10.53 7.14 5.00 28.57 15.79 9.52 5.26 32.00 9.52 11.76 15.16
Yi-VL-34B [63] 27.59 22.22 15.00 28.57 10.00 50.00 27.78 16.67 42.86 42.31 22.73 17.65 27.39
InternVL-Chat-V1-5 (26B) [64] 21.88 18.75 23.81 37.50 27.27 52.38 29.41 33.33 26.32 44.00 13.64 20.00 28.97
MiniCPM-Llama3-V [65] 37.50 16.67 33.33 71.43 41.18 47.06 38.46 16.67 41.67 57.14 33.33 25.00 38.58
Cogvlm2-llama3 [66] 22.58 22.22 27.78 30.77 15.00 61.90 35.71 45.00 14.29 51.85 28.57 38.89 33.06
Idefics2-8B [67] 23.33 27.78 23.53 20.00 13.64 50.00 40.00 22.73 11.11 41.67 14.29 40.00 26.72
Bunny-Llama-3-8B-V [3] 15.38 22.22 18.75 40.00 5.00 28.57 29.41 23.81 10.00 40.00 10.00 26.67 22.22
LLaVA-RLHF (7B) [68] 86.36 50.00 50.00 46.67 40.00 78.95 81.82 38.10 57.89 68.18 29.41 56.25 57.01
Deepseek-VL-7B-Chat [29] 30.00 20.00 27.78 43.75 31.82 71.43 77.78 45.45 47.62 57.69 14.29 38.89 42.34
Mplug-owl2-llama2-7B [69] 38.10 17.65 22.22 61.54 25.00 57.14 76.92 40.00 47.37 65.22 26.32 46.15 42.86
Phi-3-vision (4B) [70] 19.35 18.75 26.32 43.75 19.05 55.56 58.82 40.00 22.22 48.00 33.33 31.25 34.03
Bunny-MMR-3B 11.11 13.33 26.32 53.85 4.76 40.0 29.41 28.57 33.33 33.33 9.52 14.29 23.91
6 Experiment
6.1 Implementation Details
Training details.
In the visual and language alignment stage, we train only the connector for one epoch with a learning rate of 1e-5. In the fine-tuning phase, we fine-tune both the connector and the language model using LoRA [71] with a learning rate of 2e-4. The training is implemented in PyTorch using 8 Nvidia A100 GPUs in an internal server.
Evaluation benchmark.
[…]
6.2 Result Analysis of MMR Benchmark
Comparison of MLLMs on the MMR benchmark.
We evaluate 6 closed-source and 12 leading open-source models on the MMR benchmark, as shown in Tab. 2. Our findings reveal that both open-source and closed-source models suffered from the fragility of leading questions despite understanding the visual content. GPT-4o achieved the highest performance among the closed-source models, while Mini-Gemini-HD-34B led among the open-source models, achieving an RA metric of 69.0%. However, GPT-4o is more susceptible to negative questions, with an MR of 19.46%, compared to Minigemini-HD-34B (15.16% MR).
Performance of MLLMs on different types of intrusive questions.
Tab. 2 reveals that: 1) The model is particularly vulnerable to issues involving character-level details, numerical data, positional context within phrase-level categories, expertise knowledge, and relative relationships between objects within sentence-level categories. 2) Subcategories with high positive comprehension capabilities tend to exhibit greater resilience to intrusive questions. These insights highlight the imperative for a balanced enhancement of the model’s knowledge base and comprehension capabilities.
Table 3: Comparison instruction dataset on the MMR benchmark. We train with these datasets in the same backbone (SigLIP+Phi-2).
Dataset Size Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. RA $\uparrow$
LLaVA 158k [62] 158k 22.50 27.27 25.00 12.50 58.33 16.67 9.09 36.36 42.86 25.81 45.83 40.91 29.67
SVIT 158k [47] 158k 30.00 27.27 16.67 37.50 66.67 25.00 18.18 40.91 38.10 19.35 58.33 31.82 33.67
LLaVA+SVIT 316k 32.50 22.73 16.67 25.00 66.67 25.00 13.64 31.82 38.10 22.58 50.00 40.91 32.00
LRV [9] 340k 27.50 31.82 16.67 25.00 41.67 25.00 18.18 40.91 19.05 29.03 41.67 45.45 30.00
MMR-data 48k 40.00 18.18 50.00 33.33 70.83 25.00 22.73 63.64 57.14 41.94 58.33 40.91 43.33
MMR-data 112k 42.50 22.73 45.83 37.50 70.83 29.17 27.27 63.64 52.38 51.61 62.50 40.91 45.67
Dataset Size Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
LLaVA 158k [62] 158k 65.38 50.00 64.71 76.92 30.00 78.95 75.00 60.00 52.63 66.67 38.89 35.71 57.62
SVIT 158k [47] 158k 50.00 50.00 77.78 25.00 20.00 64.71 63.64 57.14 57.89 71.43 22.22 41.67 50.73
LLaVA+SVIT 316k 53.57 61.54 77.78 60.00 23.81 68.42 72.73 63.16 55.56 74.07 33.33 43.75 56.95
LRV [9] 340k 57.69 53.33 77.78 53.85 47.37 71.43 55.56 57.14 78.95 64.00 41.18 37.50 58.90
MMR-data 48k 44.83 66.67 25.00 38.46 22.73 70.00 44.44 33.33 33.33 45.83 30.00 43.75 40.91
MMR-data 112k 34.62 54.55 31.25 30.77 22.73 68.18 33.33 26.32 35.29 38.46 25.00 43.75 36.87
6.3 Result Analysis of MMR-data
Comparison of instruction tuning datasets generated by GPT.
We compare the generated dataset with previous GPT-4V generated instruction tuning datasets, such as LLaVA 158k [22], SVIT 158k [47], and LRV [9] on the proposed MMR benchmark (as shown in Tab. 3) and general benchmarks (as shown in Tab. 5). The results demonstrate that our dataset enables the model to outperform the model trained with previous datasets on both MMR and general benchmarks.
Comparison of MLLMs.
We augment our generated MMR-data with additional datasets (Sec. 4) to form a comprehensive new dataset. We validate this expanded dataset on the proposed MMR and several general benchmarks. As shown in Tab. 2 and Tab. 4, Experiments show that the 3-billion parameter model trained on this enriched dataset outperforms the LLaVA-v1.5-7B on most benchmarks. Remarkably, it also exceeds the performance of some models with 13-billion parameters on several benchmarks.
Table 4: Comparison to leading MLLMs on 10 benchmarks. §We evaluate the officially released checkpoint by ourselves.
[…]
IDEFICS-80B [75] OpenCLIP-H (1.0B) LLaMA-65B 54.6 54.5 60.0
BLIP-2 [76] EVA01-CLIP-G (1.0B) Vicuna-13B 41.0
InstructBLIP [77] EVA01-CLIP-G (1.0B) Vicuna-13B 49.5 83.7
BLIP-2 [76] EVA01-CLIP-G (1.0B) Flan-T5-XXL (11B) 1293.8 290.0 35.4 34.0 65.0 44.6
InstructBLIP [77] EVA01-CLIP-G (1.0B) Flan-T5-XXL (11B) 1212.8 291.8 35.7 33.8 47.9
Shikra-13B [78] CLIP-L (0.4B) Vicuna-13B 77.4
LLaVA-v1.5-13B (LoRA) [79] CLIP-L (0.4B) Vicuna-13B 1541.7 300.4§ 68.4§ 68.5 61.3 40.0§ 33.2§ 80.0 63.3 86.7
VILA-13B [80] CLIP-L (0.4B) LLaMA2-13B 1570.1 - 70.3 - 62.8 - - 80.8 63.3 84.2
SPHINX-Plus [49] Mixture of Visual Experts (MoV) LLaMA2-13B 1457.7 283.6 71.0 - 74.8 - - - - 89.1
LLaVA-Llama-3-8B-v1.1 [81] CLIP-L (0.4B) Llama-3-8B 1469 349 72.3 - - 36.8 - - 62.6 86.4
InstructBLIP [77] EVA01-CLIP-G (1.0B) Vicuna-7B 33.9 36.0 53.4 49.2
MiniGPT-v2 [82] EVA01-CLIP-G (1.0B) LLaMA2-7B 60.3
IDEFICS-9B [75] OpenCLIP-H (1.0B) LLaMA-7B 45.3 48.2 50.9
LLaVA-v1.5-7B (LoRA) [79] CLIP-L (0.4B) Vicuna-7B 1476.9 267.9§ 66.1§ 66.1 60.1 34.4§ 31.7§ 79.1 63.0 86.4
mPLUG-Owl2 [69] CLIP-L (0.4B) LLaMA2-7B 1450.2 313.2 66.0 66.5 57.8 32.7 32.1 79.4 56.1 85.8
VILA-7B [80] CLIP-L (0.4B) LLaMA2-7B 1533.0 - 68.9 - 61.1 - - 79.9 62.3 85.5
Shikra-7B [78] CLIP-L (0.4B) Vicuna-7B 60.2 58.8
SPHINX-Intern2 [49] Mixture of Visual Experts (MoV) InternLM2-7B 1260.4 294.6 57.9 - 68.8 - - 75.5 56.2 86.9
Mini-Gemini [23] CLIP Vicuna-7B 1523 316 - - - 36.1 32.8 65.2 - -
MM1-7B-Chat [27] ViT-H (5B) MM1-7B 1529.3 328.9 72.3 - 64.0 37.0 35.6 82.8 - 86.6
LLaVA-Phi-3-mini [81] CLIP-L (0.4B) Phi-3 (4B) 1477 313 69.2 - - 41.4 - - 61.5 87.3
Gemini Nano-2 [1] Gemini Nano-2 (3.25B) 32.6 67.5
MM1-3B-Chat [27] ViT-H (5B) MM1-3B 1482.5 279.3 67.8 - 63.0 33.9 33.7 82.0 - 87.4
MobileVLM [83] CLIP-L (0.4B) MobileLLaMA (2.7B) 1288.9 59.6 59.0 84.9
MobileVLM V2 [84] CLIP-L (0.4B) MobileLLaMA (2.7B) 1440.5 63.2 61.1 84.7
ALLaVA-Longer [46] SigLIP-SO (0.4B) Phi-2 (2.7B) - - 64.6 - 65.6 33.2 - - 50.0 -
TinyLLaVA-share-Sig-Phi [85] SigLIP-SO (0.4B) Phi-2 (2.7B) 1464.9 - 66.9 - - - - 79.9 62.0 86.4
Bunny-MMR-3B SigLIP-SO (0.4B) Phi-2 (2.7B) 1515.1 262.9 70.7 69.7 63.2 38.2 34.0 80.1 62.1 86.0
Bunny-MMR-4B SigLIP-SO (0.4B) Phi-3 (4B) 1511.2 313.2 72.3 71.8 64.1 40.9 38.8 80.4 62.2 86.3
Bunny-MMR-8B SigLIP-SO (0.4B) Llama-3-8B 1606.1 331.8 77.6 76.3 65.1 42.6 38.7 81.5 63.4 86.3
Ablation about the unpaired positive and negative samples.
To evaluate the impact of unpaired positive and negative samples, we generate negative samples and combine them with the original LLaVA samples in two ways: 1) by merging them with the original samples (LLaVA 24K merge neg.), and 2) by treating them as independent samples (LLaVA 24K + neg). Results in Tab. 6 show that the unpaired methods were significantly less effective than the paired method on both MMR and general benchmarks.
Ablation about the combination of positive and negative samples.
Table 5: Comparison between different instruction tuning data on general benchmarks. We train these datasets in the same backbone (SigLIP+Phi-2). ‘+’ denote combining two datasets.
[…]
LLaVA 158k [62] 158k 29.67 57.62 995.87 282.86 51.18 53.18 74.84
SVIT 158k [47] 158k 33.67 50.73 1190.98 269.28 53.02 56.10 75.27
LLaVA+SVIT 316k 32.00 56.95 1198.35 256.43 58.52 57.47 78.88
LRV [9] 340k 30.00 58.90 1001.79 228.57 57.51 57.04 79.68
MMR-data 48k 43.33 40.91 1321.89 255.36 63.11 63.32 79.62
Table 6: Ablation about data combination. merge: merging positive and negative conversations into one sample. pos. and neg. denote the positive and negative samples, respectively. We train these datasets in the same backbone (SigLIP+Phi-2).
[…]
LLaVA 24K 27.33 59.41 1039.62 265.36 41.48 42.53 82.51
LLaVA 24K merge neg. 21.33 63.01 1308.59 254.64 15.53 18.21 82.22
LLaVA 24K + neg. 31.33 55.45 1174.62 270.35 49.61 53.78 84.70
MMR-data (only neg.) 26.00 56.18 604.58 200.00 59.53 60.05 54.47
MMR-data (only pos.) 27.33 63.72 648.91 233.57 60.81 61.25 70.16
MMR-data (merge neg. and pos.) 27.67 63.27 908.48 221.07 59.53 59.70 79.83
MMR-data (pos. + neg.) 43.33 40.91 1321.89 255.36 63.11 63.32 79.62
Table 7: Ablation about the combination of the proposed MMR dataset with other datasets. We train these datasets with the same vision encoder (SigLIP).
[…]
Mix 106k Phi-2 44.00 43.59 1410.01 266.07 66.14 66.15 60.07 36.70 32.90 75.71 56.39 84.89
Mix 106k-replace MMR 12k Phi-2 45.33 41.63 1432.51 271.78 61.43 60.91 57.97 35.00 33.20 75.75 56.50 83.64
Mix 106k-replace MMR 24k Phi-2 47.00 40.51 1421.54 248.57 67.26 66.32 60.37 38.00 34.00 75.95 56.92 84.40
Mix 106k-replace MMR 48k Phi-2 47.33 38.26 1419.26 239.29 66.76 65.64 60.61 35.60 33.70 76.24 56.96 83.63
Mix 106k + MMR 12k Phi-2 50.00 36.71 1388.89 252.5 65.86 66.07 60.51 37.30 33.00 75.86 56.87 84.06
Mix 106k + MMR 24k Phi-2 48.67 38.66 1400.74 244.64 66.98 65.98 60.62 37.10 32.90 76.05 56.96 83.73
Mix 106k + MMR 48k Phi-2 47.00 38.16 1410.40 247.5 65.86 66.24 60.85 36.80 33.90 73.45 57.05 84.98
Base 695k Phi-2 55.33 27.19 1492.17 274.29 68.27 68.56 62.31 37.40 32.80 79.85 61.84 86.52
Base-replace MMR 112k (649k) Phi-2 57.00 22.97 1498.57 255.35 70.35 68.90 62.88 35.70 33.40 79.69 62.08 85.13
Weight merge Phi-2 58.33 23.91 1515.06 262.86 70.74 69.67 63.23 38.20 34.00 80.14 62.10 85.98
Base 695k Phi-3 54.67 29.61 1462.07 335.35 70.57 71.56 63.08 40.30 39.00 80.34 61.69 85.21
Base-replace MMR 112k (649k) Phi-3 58.33 24.57 1520.43 283.93 72.43 71.23 63.89 41.30 38.50 80.14 61.74 85.88
Weight merge Phi-3 58.33 27.39 1511.16 313.2 72.31 71.82 64.12 40.90 38.80 80.44 62.21 86.28
Base 695k Llama-3-8B 58.00 26.27 1548.53 313.57 76.63 75.34 64.62 40.70 38.60 80.99 62.82 85.30
Base-replace MMR 112k (649k) Llama-3-8B 61.00 21.12 1600.46 318.93 75.22 75.26 64.62 41.70 37.90 81.03 62.72 86.30
Weight merge Llama-3-8B 59.33 24.58 1606.12 331.79 77.63 76.29 65.11 42.60 38.70 81.52 63.44 86.25
To explore the influence of positive and negative samples on model performance, we experiment with various training configurations, including exclusive positive or negative samples, merging positive and negative conversations into one sample, and combining positive and negative samples independently. The results, outlined in Tab. 6 and Tab. 3, demonstrate a decrease in performance on the MME benchmark when using solely positive or negative samples, while directly merging the two types of data yielded suboptimal results. Surprisingly, treating positive and negative samples as separate entities led to the most favorable outcomes. These findings underscore the importance of considering the nuanced combination between positive and negative samples. Further details and analyses are available in the Appendix.
Ablation about combining the proposed MMR-data with other datasets.
To evaluate the impact of integrating the MMR-data with other datasets, we conduct experiments using an expanded dataset. We sample 24k instances from the LLaVA 158k dataset and combine them with other datasets to construct 106k samples (Mix 106k), maintaining proportional representation. We explore various combination methods, such as replacing the original LLaVA data or directly adding MMR data. Our results, presented in Tab. 7, reveal that adding new data does not consistently improve performance; replacement yields more substantial enhancements. The model achieves optimal performance when the replacement proportion closely matches the original distribution.
Ablation about upscaling the dataset.
We conduct an ablation study to examine the effects of upscaling the dataset to 695k (Base 695k) as described in Sec. 6.1. We then replace the LLaVA 158k with the MMR-data 112k. Additionally, we explore model merging by averaging the weights of two models (denoted as Merge). The results, presented in Tab. 7, demonstrate that upscaling with paired positive and negative samples significantly enhances MLLM performance. Furthermore, weight merging further boosts performance, underscoring the effectiveness of our proposed methods.
7 Conclusion and Discussion
Conclusion.
In this paper, we explored the intricate challenges of MLLMs in accurately interpreting visual content and generating precise responses to intrusive questions. We revealed that MLLMs may provide inaccurate answers despite demonstrating a nuanced understanding of visual content. To quantitatively assess the phenomenon in MLLMs, we introduced MMR, a comprehensive evaluation framework tailored to measure MLLMs’ comprehension of visual content and their resilience to negative questions. Additionally, we proposed a data construction pipeline and introduced high-quality fine-tuning data to enhance the robustness and understanding abilities of MLLMs. Our research underscores the critical need for improved evaluation methodologies and data strategies to advance MLLM capabilities in real-world scenarios.
Broader impact.
In this paper, we study how to evaluate and improve MLLMs’ robustness when answering challenging visual questions. The research will benefit the general development and applications of MLLMs.
Limitation.
Due to the cost of dataset construction, we contribute 112K instruction tuning data. Such adversarial datasets can be scaled up with more quota.
Acknowledgments and Disclosure of Funding
This work is funded by NSFC-62306046.
Appendix A MMR Benchmark Details
A.1 Category Definition and Negative Question Design
Character-Level
1. Character/Number: Recognize characters and numbers in an image, and some reasoning is based simply on the content of the characters or numbers in the image.
Adding, deleting, modifying, or swapping positions of characters or numeric content in an image.
Attribute-Level
1. Presence: Determine whether a person, animal, or object is present in the image.
Ask questions by replacing or deleting people, animals, or objects in the image.
2. Color/Texture: Determine whether the color or texture of an object in an image is correct.
Replace the color of an object in an image or replace it with an object of similar texture to ask questions.
3. Number: Count the number of people, animals, or objects in an image.
Modify the number of people, animals, or objects in the image to ask questions.
4. Shape: Determine if an object is shaped correctly.
Modify the shape of objects in the image to ask questions.
5. Posture: Determine if a person, animal, or object is in the correct posture.
Ask questions by substituting gestures of characters, animals, or objects.
6. Position: Determine whether the absolute or relative position of a person, animal, or object is correct.
Change the absolute or relative position of a character, animal, or object to ask a question.
Context-Level
1. Abstract Knowledge: Assess understanding of abstract concepts such as emotions, aesthetics, connotations, symbols, culture, news, allusions, legends, common sense, functions, etc.
No abstract knowledge is involved, and questions are asked only visually through images.
2. Concrete Knowledge: Evaluate specific factual information or well-defined objects and scenarios depicted in the image, such as landmarks, celebrities, well-known objects, etc.
Changing the attributes or names of landmarks, famous objects, or famous people.
3. Expertise: Test specialized knowledge related to specific fields or domains illustrated in the image, such as specific knowledge in various vertical fields, industries, disciplines, etc.
Ask questions with the wrong expertise.
4. Activity: Identify and interpret actions or events occurring within the image, requiring a dynamic understanding of the scene.
Ask questions with the wrong activity.
5. Relationships: Analyze and comprehend the interactions or relationships between multiple entities within the image, such as social dynamics or physical connections.ch as social dynamics or physical connections.
Change the relationship between entities in an image to ask questions.
A.2 Benchmark Data Distribution.
We have collected images that matched 12 subcategory themes from the web and some test sets, such as Flickr-30k [86], and facilitated setting misleading prompts. We meticulously design questions and standard answers based solely on the image information without relying on the original captions. The distribution of the MMR benchmark is shown in Fig. 4. It categorizes questions into three main types: "context" (120 questions), “character” (40 questions), and “attribute” (140 questions). The “context” category includes subtypes such as relationships (22), activity (24), expertise (31), concrete (21), and abstract (22) questions. The “character” category comprises character/number questions (40). The “attribute” category encompasses presence (22), color/texture (24), number (24), shape (24), posture (24), and position (22) questions. This distribution highlights a comprehensive and balanced approach, ensuring a well-rounded benchmark with detailed subcomponents across various dimensions of question types.
Figure 4: Distribution of MMR benchmark.
A.3 MLLMs Evaluation Details.
We evaluated 18 top-performing MLLMs on the MMR benchmark without adjusting parameter settings. We used the official inference code and parameter settings for each MLLM to ensure the models perform at their best during the inference process.
Figure 5: Visualization of fine-tuning loss of different data composition (version 0).
Figure 5 shows that the model’s accuracy in handling negative questions is lower than that for positive questions. If only positive questions are evaluated, as is the case with most existing benchmarks, many models achieve an accuracy rate above 70%, demonstrating good performance. However, overall, the model’s performance on negative questions is not ideal, indicating that existing models still struggle with leading questions. The higher accuracy on positive questions shows that the model can understand the content of images. In comparison, the lower accuracy on negative questions suggests that the model tends to overlook visual content when faced with leading questions, resulting in incorrect answers. This highlights the necessity of establishing the MMR Benchmark to evaluate the robustness of models against leading questions quantitatively.
Is the proposed MMR benchmark effective for MLLMs?
MMStar [38] critically examines prevailing evaluation methods and brings to light a significant issue: many instances don’t require visual stimuli for accurate responses. Answers can be inferred directly from the questions, choices, or the intrinsic world knowledge within LLMs. This trend pervades existing benchmarks. To investigate if this issue persists in our benchmark, we conduct experiments using both image and text inputs, as well as only text inputs. The results, showcased in Tab. 8 and Tab. 9, reveal that without text, the model’s performance on our benchmark dips below random selection. This underscores the necessity of multimodal input for our proposed benchmark and highlights that our benchmark is not solely reasoned out by text.
Table 8: Performance of MLLMs with images as input on the MMR benchmark. §We evaluate the officially released checkpoint by ourselves. Abbreviations: Char/Num. (Character/Number), Pres. (Presence), Color/Tex. (Color/Texture), Num. (Number), Shape (Shape), Posture (Posture), Pos. (Position), Abstract. (Abstract Knowledge), Concrete. (Concrete Knowledge), Expert (Expertise), Act. (Activity), Rel. (Relationships). $\uparrow$: higher is better, $\downarrow$: lower is better.
[…]
Random 25 25 25 25 25 25 25 25 25 25 25 25 25
Phi-3-vision (4B) [70] 62.50 59.09 58.33 37.50 70.83 33.33 31.82 54.55 66.67 41.94 58.33 50.00 52.33
Bunny-Llama-3-8B-V [3] 55.00 63.64 54.17 37.50 79.17 62.50 54.55 72.73 85.71 48.39 75.00 50.00 60.67
Idefics2-8B [67] 57.50 59.09 54.17 50.00 79.17 41.67 27.27 77.27 76.19 45.16 75.00 40.91 56.67
Method Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
Random 50 50 50 50 50 50 50 50 50 50 50 50 50
Phi-3-vision (4B) [70] 19.35 18.75 26.32 43.75 19.05 55.56 58.82 40.00 22.22 48.00 33.33 31.25 34.03
Bunny-Llama-3-8B-V [3] 15.38 22.22 18.75 40.00 5.00 28.57 29.41 23.81 10.00 40.00 10.00 26.67 22.22
Idefics2-8B [67] 23.33 27.78 23.53 20.00 13.64 50.00 40.00 22.73 11.11 41.67 14.29 40.00 26.72
Table 9: Performance of MLLMs without images as input on the MMR benchmark. §We evaluate the officially released checkpoint by ourselves. Abbreviations: Char/Num. (Character/Number), Pres. (Presence), Color/Tex. (Color/Texture), Num. (Number), Shape (Shape), Posture (Posture), Pos. (Position), Abstract. (Abstract Knowledge), Concrete. (Concrete Knowledge), Expert (Expertise), Act. (Activity), Rel. (Relationships). $\uparrow$: higher is better, $\downarrow$: lower is better.
[…]
Random 25 25 25 25 25 25 25 25 25 25 25 25 25
Phi-3-vision (4B) [70] 15.0 13.64 20.83 0.0 12.5 8.33 9.09 27.27 19.05 16.13 12.5 13.64 14.00
Bunny-Llama-3-8B-V [3] 22.5 13.64 16.67 4.17 4.17 16.67 22.73 22.73 28.57 19.35 8.33 27.27 17.33
Idefics2-8B [67] 15.0 4.55 8.33 4.17 8.33 4.17 0.0 31.82 23.81 3.23 0.0 13.64 9.76
Method Char/Num Pres. Color/Tex Num. Shape Posture Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
Random 50 50 50 50 50 50 50 50 50 50 50 50 50
Phi-3-vision (4B) [70] 72.73 72.73 58.33 100.0 66.67 75.0 81.82 57.14 60.0 72.22 62.5 72.73 69.34
Bunny-Llama-3-8B-V [3] 64.0 75.0 63.64 83.33 91.67 69.23 44.44 64.29 50.0 70.0 80.0 57.14 67.09
Idefics2-8B [67] 77.78 90.91 80.0 85.71 85.71 92.31 100.0 56.25 61.54 95.24 100.0 80.0 82.10
Appendix B MMR Data Collection and Preliminary Experiment Details
B.1 Training Details
We employed a two-stage training strategy with specific hyperparameters for pre-training and fine-tuning, as detailed in Tab. 10. For testing, we used greedy search to ensure reproducibility.
Table 10: Hyperparameters.
batch size lr lr schedule lr warmup ratio weight decay epoch optimizer DeepSpeed stage
Pretrain 256 1e-5 cosine decay 0.03 0 1 AdamW 2
Finetune 128 2e-4 cosine decay 0.03 0 1 AdamW 3
B.2 Can we simply introduce negative samples to train MLLMs?
We observed that many MLLMs could understand the content but struggled to answer questions. We hypothesized that the insufficiency of negative samples might be the cause. To address this, we generated additional negative samples and incorporated them into the training process. We sampled 7k and 24k samples from LLaVA and used GPT-4V to generate the negative samples.
Prompt. We employ two types of prompts for this task version 0, which restricts only the number of problems, and version 1, which restricts both the type and number of problems. Their common prompt is partially displayed in Fig. 6, while their respective different prompts are displayed in Fig. 7 and Fig. 8, respectively.
STEP-1: Parse and extract information based on the Input JSON and Information Topics: 1. Determine if there is any very clear text or numbers (You must make sure the identification is right) in the image. If yes, extract the corresponding text or numbers along with its position. 2. Determine if there is any object in the image, including the properties of that object, such as name/category/color/texture/shape/pose (The properties of an object may not be unique, such as color, it may be a mixture of different colors), position, and quantity (You must make sure the identification is right). 3. Determine if there is any person in the image, including attributes such as name, position, gender, and facial expression (You must make sure the identification is right. You must differentiate between a portrait in a photo or poster and a real person, don’t confuse them.). 4. Understand the events occurring in the image at a global level, and assess whether they relate to culture, emotions, or knowledge (You must make sure the identification is right). 5. Understand the relationships between objects within the image, such as relative positions or hierarchical relationships (You must make sure the identification is right). 6. Understand the relationships between people and objects, such as someone riding a bike or a person’s actions (You must make sure the identification is right). 7. If there are two or more people present, extract activities they are engaged in.
Figure 6: Common Prompt for information extraction.
STEP-2: Generate questions and corresponding correct answers based on the extracted information in STEP-1, Input Image, and Question Types (Do not output):
Question Types:
1. If there is no text/number recognized or you do not make sure the text/number is correct, do not ask this type of question. If text is present, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. 2. If characters or numbers are present in the image, modify one letter of a word or one digit of a number, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 3. Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. 4. Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 5. If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 6. Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. 7. Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 8. Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). 9. Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
Figure 7: GPT4 prompt for version 0.
STEP-2: Generate questions and corresponding correct answers based on the extracted information in STEP-1, Input Image, and Question Types (Do not output):
Question Types:
1. Character: (a) If no text/number is recognized or the text/number recognized is not sure, do not ask questions about text/number. If text exists, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. (b) If there are characters or numbers in the image, modify one letter or one digit, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 2. Semantic: (a) Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. (b) Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 3. Understanding: (a) Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. (b) Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). (d) Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 2 character-type questions, 4 semantic-type questions, and 2 understanding-type questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
Figure 8: GPT4 Prompt for version 1.
B.2.1 Training and results of the version 0.
We train MLLMs using the sampled data and synthetically generated data. The hyper-parameters used for training remain consistent with those outlined in Tab. 10. We first explore the model’s performance using different data ratios: 1) sampled data alone, 2) sampled data concatenated with generated data, and 3) sampled data combination with generated data. The loss curves for the fine-tuning process are illustrated in Fig. 9. The loss curves in Fig. 9 reveal that incorporating generated negative data, whether through concatenation or combination, leads to a further reduction in model loss compared to training solely on sampled data. Notably, the lowest loss is achieved when utilizing a combination of both sampled and generated data.
Tab. 11 and Tab. 12 present the model’s performance on MMR and other benchmarks across the different data compositions. As expected, training exclusively on sampled data yields the lowest overall performance. Interestingly, while splicing sampled data with generated negative samples leads to a performance decrease on the generic benchmark, it demonstrates improvements on certain phantom benchmarks. The most effective approach proved to be combining the original sampled data with generated negative samples. This strategy consistently enhances performance across most benchmarks, often by a significant margin.
Figure 9: Visualization of fine-tuning loss of different data composition (version 0).
Table 11: Ablation about only introducing negative samples on the MMR benchmark. Concat. and Neg. denotes concatenate and negative samples, respectively.
[…]
LLaVA 7K 2.50 0.00 4.17 0.00 0.00 0.00 4.55 0.00 4.76 0.00 0.00 4.55 1.67
LLaVA 7K Concat. Neg. 2.50 0.00 4.17 4.17 20.83 0.00 4.55 0.00 0.00 3.23 4.17 4.55 4.00
LLaVA 7K Conbine Neg. 5.00 4.55 8.33 16.67 25.00 4.17 9.09 18.18 4.76 3.23 16.67 9.09 10.00
LLaVA 24K 17.50 22.73 20.83 12.50 45.83 16.67 13.64 50.00 28.57 29.03 45.83 31.82 27.33
LLaVA 24K Concat. Neg. 20.00 13.64 8.33 16.67 58.33 12.50 4.55 40.91 19.05 9.68 25.00 31.82 21.33
LLaVA 24K Conbine Neg. 27.50 27.27 25.00 20.83 54.17 16.67 9.09 45.45 38.10 25.81 54.17 36.36 31.33
Method Char/Num Pres. Color/Tex Num. Shape Stance Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
LLaVA 7K 83.33 100.00 80.00 100.00 100.00 100.00 75.00 100.00 80.00 100.00 100.00 80.00 93.06
LLaVA 7K Concat. Neg. 88.89 100.00 66.67 87.50 44.44 100.00 66.67 100.00 100.00 92.31 75.00 75.00 84.42
LLaVA 7K Conbine Neg. 86.67 83.33 71.43 66.67 57.14 92.31 66.67 66.67 92.31 93.75 55.56 81.82 77.61
LLaVA 24K 72.00 54.55 68.75 75.00 38.89 77.78 62.50 45.00 68.42 62.50 38.89 46.15 59.41
LLaVA 24K Concat. Neg. 52.94 62.50 85.71 63.64 26.32 80.00 87.50 50.00 77.78 83.33 53.85 50.00 63.01
LLaVA 24K Conbine Neg. 59.26 50.00 60.00 58.33 31.58 78.95 77.78 47.37 60.00 68.00 35.00 42.86 55.45
Table 12: Ablation about only introducing negative samples on general benchmarks. Concat. and Neg. denotes concatenate and negative samples, respectively.
[…]
LLaVA 7K 719.07 272.86 1.68 2.23 74.70
LLaVA 7K Concat. Neg. 1095.19 267.5 0.50 1.03 82.69
LLaVA 7K Conbine Neg. 883.14 287.14 4.76 4.90 82.99
LLaVA 24K 1039.62 265.36 41.48 42.53 82.51
LLaVA 24K Concat. Neg. 1308.59 254.64 15.53 18.21 82.22
LLaVA 24K Conbine Neg. 1174.62 270.35 49.61 53.78 84.70
Ablation about conversation rounds. Our initial concatenation approach directly appended all conversation rounds of a negative sample to the original data. However, this might lead to performance degradation due to data imbalance. To investigate this, we conducted an ablation study on the 24k dataset, focusing on the number of conversation rounds included from negative samples. Fig. 10 presents the model’s training loss across different numbers of conversation rounds. Additionally, Tab. 14 and Tab. 13 showcase the corresponding performance on the MMR and generalized benchmarks, respectively. The experimental results indicate that using fewer concatenated conversation rounds does not significantly degrade model performance. However, a minor performance drop is still observed. We hypothesize that this drop stems from the increased variability within each training sample when fewer rounds are used. This variability might make it more challenging for the model to learn consistent patterns and generalize effectively.
Figure 10: Visualization of fine-tuning loss of different conversation rounds.
Table 13: Ablation about only introducing negative samples on general benchmarks. Concat. and Neg. denotes concatenate and negative samples, respectively.
[…]
LLaVA 24K 1039.62 265.36 41.48 42.53 82.51
LLaVA 24K Concat. Neg. (r1) 1304.77 268.21 40.87 41.87 81.65
LLaVA 24K Concat. Neg. (r2) 1336.41 258.57 24.74 18.21 80.77
LLaVA 24K Concat. Neg. (r4) 1304.25 259.29 25.00 28.52 82.42
LLaVA 24K Concat. Neg. (all ) 1308.59 254.64 15.53 18.21 82.22
Table 14: Ablation about only introducing negative samples on the MMR benchmark. Concat. and Neg. denotes concatenate and negative samples, respectively.
[…]
LLaVA 24K 17.50 22.73 20.83 12.50 45.83 16.67 13.64 50.00 28.57 29.03 45.83 31.82 27.33
LLaVA 24K Concat. Neg. (r1) 22.50 22.73 16.67 16.67 45.83 20.83 9.09 45.45 23.81 22.58 45.83 27.27 26.33
LLaVA 24K Concat. Neg. (r2) 17.50 13.64 8.33 12.50 58.33 16.67 9.09 45.45 38.10 16.13 37.50 31.82 24.67
LLaVA 24K Concat. Neg. (r4) 17.50 9.09 20.83 16.67 70.83 16.67 13.64 45.45 33.33 19.35 45.83 36.36 28.00
LLaVA 24K Concat. Neg. (all) 20.00 13.64 8.33 16.67 58.33 12.50 4.55 40.91 19.05 9.68 25.00 31.82 21.33
Method Char/Num Pres. Color/Tex Num. Shape Stance Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
LLaVA 24K 72.00 54.55 68.75 75.00 38.89 77.78 62.50 45.00 68.42 62.50 38.89 46.15 59.41
LLaVA 24K Concat. Neg. (r1) 65.38 61.54 73.33 69.23 42.11 73.68 80.00 47.37 73.68 72.00 38.89 53.85 62.20
LLaVA 24K Concat. Neg. (r2) 65.00 72.73 85.71 75.00 33.33 76.47 80.00 47.37 60.00 77.27 43.75 53.33 62.44
LLaVA 24K Concat. Neg. (r4) 63.16 80.00 66.67 60.00 19.05 75.00 66.67 47.37 65.00 71.43 42.11 42.86 56.48
LLaVA 24K Concat. Neg. (all) 52.94 62.50 85.71 63.64 26.32 80.00 87.50 50.00 77.78 83.33 53.85 50.00 63.01
B.2.2 Training and results of the version 1.
To further investigate the model’s performance under constraints on question types, we introduce a modified prompt and generate a new 24k dataset for experiments. Interestingly, despite prompting for negative samples, GPT-4V still produced a mix of positive and negative examples, with an approximate ratio of 1:13. Utilizing only this generated dataset, we have conducted experiments to analyze the impact of different problem types. Fig. 11 illustrates the training loss, while Tab. 15 and Tab. 16 present the corresponding performance results on benchmarks.
Figure 11: Visualization of fine-tuning loss of different data composition (version 1).
Table 15: Ablation about different question types.
[…]
Character 869.21 203.93 58.63 58.59 66.18
Semantic 1167.73 216.78 61.43 61.52 71.33
Understanding 918.38 219.64 60.15 59.88 67.93
Character + Semantic 963.81 225.36 59.98 60.05 68.84
Semantic + Understanding 1260.70 246.43 62.05 61.77 71.82
Character + Semantic + Understanding 1051.6 242.5 61.27 61.00 71.63
Table 16: Ablation about only introducing negative samples on the MMR benchmark. Concat. and Neg. denotes concatenate and negative samples, respectively.
Method Char/Num Pres. Color/Tex Num. Shape Stance Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MA $\uparrow$
Character 35.00 40.91 20.83 20.83 62.50 33.33 22.73 59.09 33.33 25.81 50.00 31.82 36.00
Semantic 27.50 31.82 33.33 33.33 79.17 29.17 18.18 45.45 47.62 35.48 58.33 45.45 39.67
Understanding 47.50 40.91 29.17 33.33 66.67 41.67 31.82 54.55 47.62 38.71 58.33 40.91 44.33
Character + Semantic 45.00 40.91 29.17 37.50 66.67 33.33 27.27 59.09 47.62 35.48 54.17 40.91 43.00
Semantic + Understanding 32.50 45.45 33.33 41.67 70.83 20.83 18.18 68.18 28.57 38.71 58.33 45.45 41.33
Character + Semantic + Understanding 42.50 36.36 29.17 37.50 70.83 29.17 27.27 54.55 42.86 35.48 58.33 45.45 42.33
Method Char/Num Pres. Color/Tex Num. Shape Stance Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
Character 46.15 35.71 61.54 66.67 21.05 60.00 44.44 38.10 58.82 68.00 29.41 50.00 48.57
Semantic 62.07 30.00 52.94 38.46 5.00 63.16 66.67 52.38 47.37 57.69 26.32 41.18 46.40
Understanding 26.92 30.77 53.33 33.33 20.00 54.55 30.00 42.86 44.44 53.85 26.32 43.75 38.99
Character + Semantic 30.77 35.71 58.82 25.00 23.81 57.89 45.45 38.10 47.37 57.69 27.78 47.06 41.63
Semantic + Understanding 51.85 23.08 52.94 28.57 15.00 72.22 63.64 31.82 68.42 55.56 30.00 44.44 45.13
Character + Semantic + Understanding 37.04 33.33 56.25 25.00 19.05 63.16 50.00 45.45 50.00 54.17 26.32 44.44 42.27
This experiment yielded several key insights: 1) Avoid rigidly specifying the number of question types, particularly for characteristics like characters. Enforcing such constraints, especially when not all images contain those characteristics, can lead to generating nonsensical question-answer pairs and consequently, performance degradation. 2) Training solely on generated data can be surprisingly effective, but only when both positive and negative sample pairs are present. This finding highlights the potential of synthetic data for training MLLMs. Notably, using only negative samples restricts the model from producing negative answers, emphasizing the importance of balanced datasets. Consequently, we shifted our focus to generating paired positive and negative samples to investigate model performance under this training paradigm further.
B.3 What constitutes an effective pair of positive and negative samples for training?
Having established the benefit of incorporating both positive and negative samples during training, we now delve into effective construction methods for such pairs. We investigate two distinct approaches: 1) Explicit Positive/negative Prompting. This method involves directly prompting the model to generate both positive and negative samples, with answers explicitly labeled as "yes" or "no." This approach offers straightforward control over the positive/negative balance in the training data (version 2). Confounding Answer Construction. This approach focuses on crafting questions that naturally lend themselves to multiple plausible answers, without explicitly forcing a "yes/no" response. This strategy aims to create a more nuanced and challenging training dataset, potentially enhancing the model’s ability to discern subtle differences in meaning and context (version 3). The GPT-4V prompts for version 2 and version 3 are shown in Fig. 12 and Fig. 13, respectively. To evaluate the effectiveness of these two versions of prompts, we generated 24k positive-negative sample pairs using each version’s prompt. We then incorporated these datasets into our training pipeline, maintaining consistent experimental settings with previous trials. Fig. 14 presents the training loss curves observed for both approaches, while Tab. 17 and Tab. 18 detail the corresponding performance results across various benchmarks. Our experimental results demonstrate that utilizing prompt version 3 to generate paired positive and negative samples yielded the strongest overall performance. Consequently, we adopted version 3 as our preferred method for generating training data in the main experiments.
STEP-2: Generate questions and corresponding correct answers based on the extracted information in STEP-1, Input Image, and Question Types (Do not output):
Question Types:
1. Ask if a certain object exists, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the object mentioned in the ques-pos and ans-pos with the visually similar nonexistent object. Ask if the model identifies the replaced object, the question is defined as ques-neg. Rhetorical questions are preferred. Provide the correct answer to ques-neg as ans-neg. 2. Ask a question about the name/category/color/texture/shape/pose of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the name/category/color/texture/shape/pose of the object mentioned in the ques-pos. Ask a question about if the modified name/category/color/texture/shape/pose of an object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 3. Ask a question about the number/position of an object in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the number/position of the object mentioned in the ques-pos with a similar number/position as interference. Ask a question about if the modified number/position of the object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 4. Ask a question about the topic of abstract-knowledge/concrete-knowledge/professional-knowledge in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the knowledge related to the topic mentioned in the ques-pos with similar knowledge as interference. Ask a distractor question about if the modified knowledge related to the topic is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. - Candidate topics of abstract knowledge: emotions, aesthetics, connotations, symbols, culture, news, allusions, legends, common sense, functions, and so on. - Candidate topics of concrete knowledge: landmarks, celebrities, well-known objects, and so on. - Candidate topics of professional knowledge: specific knowledge in various vertical fields, industries, disciplines, and so on. 5. Ask a question about the activity of an object and/or interaction/relationship between objects or persons in the image, the question is defined as ques-pos. Provide the correct answer to ques-pos as ans-pos. (2) Replace the activity/interaction/relationship mentioned in the ques-pos with a similar activity/interaction/relationship as interference. Ask a distractor question about if the modified activity/interaction/relationship is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. Provide the correct answer to ques-neg as ans-neg. 6. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. 7. NOTE: Pose questions that differ slightly from true understanding but are very similar, or add a distractor sentence to questions, so that the model is guided to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely, and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’id’: ’json[’id’]’, ’image’: ’json[’image’]’, ’conversations-pos’: [’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’,…, ’from’: ’human’, ’value’: ’ques-pos’, ’from’: ’gpt’, ’value’: ’ans-pos’], ’conversations-neg’: [’from’: ’human’, ’value’: ’ques-neg ’, ’from’: ’gpt’, ’value’: ’ans-neg’,…, ’from’: ’human’, ’value’: ’ques-neg’, ’from’: ’gpt’, ’value’: ’ans-neg’] STEP-3: Output JSON sample
Figure 12: GPT4 prompt for version 2.
STEP-2: Generate questions and corresponding correct answers based on the extracted information in STEP-1, Input Image, and Question-Answer Types: Question-Answer Types: 1. Types-1: (1) Ask if a certain object exists, the question is defined as ques-pos. (2) Operation-pos (3) Replace the object mentioned in the ques-pos with the visually similar nonexistent object. Ask if the model identifies the replaced object, the question is defined as ques-neg. Rhetorical questions are preferred. (4) Operation-neg 2. Types-2: (1) Ask a question about the name/category/color/texture/shape/pose of an object in the image, the question is defined as ques-pos. (2) Operation-pos (3) Replace the name/category/color/texture/shape/pose of the object mentioned in the ques-pos. Ask a question about whether the modified name/category/color/texture/shape/pose of an object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. (4) Operation-neg 3. Types-3:(1) Ask a question about the number/position of an object in the image, the question is defined as ques-pos. (2) Operation-pos (3) Replace the number/position of the object mentioned in the ques-pos with a similar number/position as interference. Ask a question about if the modified number/position of the object is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. (4) Operation-neg 4. Types-4:(1) Ask a question about the topic of abstract-knowledge/concrete-knowledge/professional-knowledge in the image, the question is defined as ques-pos. (2) Operation-pos (3) Replace the knowledge related to the topic mentioned in the ques-pos with similar knowledge as interference. Ask a distractor question about if the modified knowledge related to the topic is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. (4) Operation-neg Candidate topics of abstract knowledge: emotions, aesthetics, connotations, symbols, culture, news, allusions, legends, common sense, functions, and so on. Candidate topics of concrete knowledge: landmarks, celebrities, well-known objects, and so on. Candidate topics of professional knowledge: specific knowledge in various vertical fields, industries, disciplines, and so on. 5. Types-5:(1) Ask a question about the activity of an object and/or interaction/relationship between objects or persons in the image, the question is defined as ques-pos. (2) Operation-pos (3) Replace the activity/interaction/relationship mentioned in the ques-pos with a similar activity/interaction/relationship as interference. Ask a distractor question about if the modified activity/interaction/relationship is correct. Rhetorical questions are preferred. The new question is defined as ques-neg. (4) Operation-neg Notes: All (2) Operation-pos is the same type: Design and provide four options based on ques-pos, including opt-pos-0: right answer with the correct reason, opt-pos-1: right answer with an incorrect reason, opt-pos-2: wrong answer with the correct reason, opt-pos-3: wrong answer with an incorrect reason. The opt-pos-0 is defined as ans-pos. All (4) Operation-neg is the same type: Design and provide four options based on ques-neg, including opt-neg-0: right answer with the correct reason, opt-neg-1: right answer with an incorrect reason, opt-neg-2: wrong answer with the correct reason, opt-neg-3: wrong answer with an incorrect reason. The opt-neg-0 is defined as ans-neg. The correct reason is the only one, while the incorrect reason can stem from various distortions and fabrications of facts, etc. Pose questions that differ slightly from true understanding but are very similar or add a distractor sentence to questions so that the model is guided to make incorrect inferences or answers. For ques-neg, rhetorical questions are preferred. STEP-3: Output JSON sample
Figure 13: Instructions for generating questions and answers based on extracted image information. GPT4 prompt for version 3.
Figure 14: Visualization of fine-tuning loss of different data compositions (version 2 and version 3).
Table 17: Ablation about different combinations of data. "Version 2/3-Pos." denotes positive samples generated by the respective prompt version, while "Version 2/3-Neg." represents the corresponding negative samples.
[…]
Version 2-Pos. 714.63 241.07 59.52 60.65 74.85
Version 2-Neg. 500.0 200.0 - - -
Version 2-Pos.+Neg. 1289.56 276.43 59.70 61.43 75.78
Version 3-Pos. 648.91 233.57 60.81 61.25 70.16
Version 3-Neg. 604.58 200.0 59.53 60.05 54.47
Version 3-Pos.+Neg. 1321.89 255.36 63.11 63.32 79.62
Table 18: Ablation about different combinations of data on MMR benchmark. "Version 2/3-Pos." denotes positive samples generated by the respective prompt version, while "Version 2/3-Neg." represents the corresponding negative samples.
[…]
Version 2-Pos. 22.5 18.18 20.83 33.33 70.83 20.83 18.18 50.0 28.57 19.35 50.0 18.18 30.33
Version 2-Neg. 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Version 2-Pos.+Neg. 20.0 9.09 20.83 33.33 58.33 25.0 13.64 45.45 23.81 19.35 50.0 18.18 27.67
Version 3-Pos. 20.0 18.18 16.67 16.67 58.33 20.83 18.18 45.45 23.81 19.35 50.0 27.27 27.33
Version 3-Neg. 27.5 9.09 16.67 8.33 45.83 20.83 0.0 45.45 23.81 32.26 45.83 31.82 26.0
Version 3-Pos.+Neg. 40.0 18.18 50.0 33.33 70.83 25.0 22.73 63.64 57.14 41.94 58.33 40.91 43.33
Method Char/Num Pres. Color/Tex Num. Shape Stance Pos. Abstract. Concrete. Expert. Act. Rel. Avg. MR $\downarrow$
Version 2-Pos. 67.86 73.33 73.68 46.67 19.05 77.27 69.23 50.0 68.42 76.0 36.84 73.33 60.94
Version 2-Neg. 100.0 100.0 100.0 0 100.0 100.0 0 100.0 100.0 100.0 100.0 100.0 100.0
Version 2-Pos.+Neg. 72.41 80.0 70.59 50.0 36.36 71.43 66.67 52.38 73.68 76.0 42.86 75.0 63.27
Version 3-Pos. 70.37 69.23 75.0 71.43 33.33 77.27 66.67 52.38 75.0 75.0 40.0 62.5 63.72
Version 3-Neg. 50.0 75.0 69.23 80.0 38.89 66.67 100.0 52.38 68.75 52.38 35.29 36.36 56.18
Version 3-Pos.+Neg. 44.83 66.67 25.0 38.46 22.73 70.0 44.44 33.33 33.33 45.83 30.0 43.75 40.91
Appendix C Examples of paired positive and negative samples generated.
Refer to caption Conversation Question: Are the people preparing for a party? Answer: No, they are preparing a meal. Question: Is there a green chair visible? Answer: No, no green chair. Question: Are there three people in the kitchen? Answer: No, only two people. Question: Is the kitchen light pink? Answer: No, it is not pink. Question: Does the window show a garden outside? Answer: No visible garden.
Figure 15: Generated conversation example by prompt version 0.
Conversation (Question type: Character) Question: Is there a number 358 visible anywhere? Answer: No, there are no numbers visible. Question: Does the text ’Happy Birthday’ appear in the kitchen? Answer: No, there is no text visible. Conversation (Question type: Semantic) Question: Is there a microwave in the image? Answer: No, there is no microwave visible. Question: Can you see a refrigerator in the scene? Answer: No, there is no refrigerator visible. Question: Is there a blender on the counter? Answer: No, there is no blender shown. Question: Are there five plates in the picture? Answer: No, there are more than five plates visible. Conversation (Question type: Understanding) Question: Are the individuals preparing for a meal? Answer: Yes, serving food. Question: Is the setting a restaurant kitchen? Answer: No, it’s a home kitchen.
Figure 16: Generated conversation example by prompt version 1.
Positive Conversation Question: Is there a bowl of food on the counter? Answer: Yes, there are multiple bowls of food on the counter. Question: What is the man wearing? Answer: The man is wearing a plaid shirt. Question: How many people are in the kitchen? Answer: There are two people in the kitchen. Question: What are the people doing? Answer: They are preparing a meal. Question: Where is the sink located in the kitchen? Answer: The sink is located towards the right. Negative Conversation Question: Is there a vase of flowers on the counter? Answer: No, there is no vase of flowers on the counter. Question: What is the man wearing? Answer: The man is wearing a green hoodie. Question: How many children are in the kitchen? Answer: There are no children in the kitchen. Question: What are some people doing? Answer: They are reading books. Question: Where is the dishwasher located in the kitchen? Answer: The dishwasher is located towards the right.
Figure 17: Generated conversation example by prompt version 2.
Positive Conversation Question: Are there two people in the kitchen? Option: {"Yes, there are two people.", "Yes, there are three people.", "No, there is only one person.", "No, there are four people."} Answer: Yes, there are two people. Question: Are the people in the kitchen cooking? Option: {"Yes, they are cooking.", "No, they are cleaning.", "No, they are eating.", "Yes, they are fixing something."} Answer: Yes, they are cooking. Question: How many bowls are there on the counter? Option: {"One bowl.", "Four bowls.", "Six bowls.", "Two bowls."} Answer: Four bowls. Question: What is the main activity happening? Option: {"Preparing a meal.", "Washing dishes.", "Playing games.", "Reading books."} Answer: Preparing a meal. Question: Is there a dishwasher in the image? Option: {"Yes, the dishwasher is there.", "No, there’s no dishwasher.", "It’s unclear.", "There is a microwave instead."} Answer: Yes, the dishwasher is there. Question: Is there a pot on the table? Option: {"Yes, there is a pot.", "No, there isn’t a pot.", "Yes, there are two pots.", "No, there are only pans."} Answer: Yes, there is a pot. Question: Are the people preparing food in a modern kitchen? Option: {"Yes, they are.", "No, it’s an old kitchen.", "Yes, in a living room.", "No, it’s a commercial kitchen."} Answer: Yes, they are.
Figure 18: Generated conversation example by prompt version 3 (part 1).
Negative Conversation Question: Are there three people in the kitchen? Option: {"Yes, there are two people.", "Yes, there are three people.", "No, only one person.", "No, four people."} Answer: No, only one person. Question: Are the people in the kitchen cleaning? Option: {"Yes, they are cooking.", "No, they are cleaning.", "No, they are eating.", "Yes, fixing something."} Answer: Yes, they are cooking. Question: How many bowls are there on the countertop? Option: {"One bowl.", "Four bowls.", "Six bowls.", "Two bowls."} Answer: Four bowls. Question: What is the secondary activity happening? Option: {"Preparing a meal.", "Washing dishes.", "Playing games.", "Reading books."} Answer: Preparing a meal. Question: Is there a microwave in the image? Option: {"Yes, the microwave is there.", "No, there’s no microwave.", "It’s unclear.", "There is a dishwasher."} Answer: There is a dishwasher. Question: Is there a pan on the table? Option: {"Yes, there is a pot.", "No, there isn’t a pot.", "Yes, there are two pans.", "No, only bowls."} Answer: Yes, there is a pot. Question: Are the people preparing food in a commercial kitchen? Option: {"Yes, they are.", "No, it’s old kitchen.", "Yes, a living room.", "No, modern kitchen."} Answer: No, modern kitchen.
Figure 19: Generated conversation example by prompt version 3 (part 2).
Appendix D Licensing and Hosting
Author Statement.
We bear all responsibilities for the licensing, distribution, and maintenance of our dataset.
License.
MMR benchmark and MMR-data are under CC BY 4.0 license.
Hosting.
MMR benchmark and MMR-data can be viewed and downloaded on GitHub at https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark or on Huggingface at https://huggingface.co/datasets/BAAI/Multimodal-Robustness-Benchmark. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in JSON file format.
Metadata.
Appendix E Datasheet and Maintenance Plan
E.1 Motivation
For what purpose was the dataset created?
Answer: The MMR benchmark is designed to evaluate MLLMs’ comprehension of visual content and robustness against misleading questions, ensuring models truly leverage multimodal inputs rather than relying solely on textual reasoning. The MMR-data is generated to enhance MLLMs’ understanding capability and robustness by providing a training set with paired positive and negative visual question-answer samples.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
Answer: The MMR benchmark and MMR-data are created by Yexin Liu (BAAI, HKUST(GZ)), Zhengyang Liang (BAAI, BUPT), Yueze Wang (BAAI), Muyang He (PKU), Jian Li (Youtu Lab, Tencent), Bo Zhao (BAAI)
Who funded the creation of the dataset?
Answer: 1) Beijing Academy of Artificial Intelligence. 2) NSFC-62306046.
E.2 Composition
What do the instances that comprise the dataset represent? (e.g., documents, photos, people, countries)
Answer: Each instance in the MMR benchmark represents an image, a question, 4 options, and a standard answer. Each instance in the MMR-data includes an image, a question, and an answer. Images are stored in JPG or PNG file format, while the questions, options, and standard answers are all stored in JSON format files.
How many instances are there in total (of each type, if appropriate)?
Answer: For the MMR benchmark, we collect 600 instances, consisting of 300 positive and 300 negative questions. The specific distribution of question types is detailed in Fig. 4 in the appendix. For MMR-data, we generate 112k samples (56k positive and 56k negative) using GPT-4o, with an average of 6.82 rounds of conversation per sample.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
Answer: All instances in the MMR benchmark and MMR-data were newly annotated by us or generated by GPT-4o.
Is there a label or target associated with each instance?
Answer: Yes, each instance provides the correct answer with an explanation.
Is any information missing from individual instances?
Answer: All instances are complete.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
Answer: For the MMR benchmark, after determining the question type, we then look for diverse instances. Each image is independent and not explicitly connected to others and two questions (positive and leading questions) are annotated with wwith each image. For MMR-data, some instances may have the same image but different questions and answers. Each instance clearly indicates the corresponding image, and each image has a unique identifier.
Are there recommended data splits (e.g., training, development/validation, testing)?
Answer: Yes, the MMR benchmark is specifically designed for evaluation, while the MMR-data is generated for training.
Are there any errors, sources of noise, or redundancies in the dataset?
Answer: No.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
Answer: The annotations of the MMR benchmark are manually annotated by us, whereas the annotations of MMR-data are generated by GPT-4o. All metadata will be publicly accessible in the dataset repository. The images are collected from websites or open-source datasets. We have taken measures such as changing the resolution, modifying the aspect ratio, changing the format, and editing to minimize the impact on the original work rights. However, it is still possible that the copyright holder may request the deletion of certain data. If this happens, we will edit the content without affecting the questions and answers. If retaining images is also not allowed, we will still keep the annotation data and provide metadata (including URL) for the corresponding images.
Does the dataset contain data that might be considered confidential?
Answer: No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
Answer: No.
E.3 Collection Process
The data collection process for the MMR benchmark is detailed in Appendix A. 2 of the main paper. When labeling the MMR benchmark, those involved in the labeling search for suitable images based on pre-defined question types, type definitions, questions, and answer requirements, and then label them. The data collection process for MMR-data is outlined in Sec. 4 of the main paper.
E.4 Uses
Has the dataset been used for any tasks already?
Answer: Yes, the MMR benchmark has been used to evaluate the understanding capabilities and robustness to leading questions in 18 different MLLMs. The MMR-data has been used to train three different MLLMs, validating the effectiveness of the dataset.
What (other) tasks could the dataset be used for?
Answer: The MMR benchmark is primarily used to comprehensively evaluate the understanding capabilities and robustness to leading questions in MLLMs. Given its inclusion of multiple different tasks, it can also be used to individually assess the ability of MLLMs on specific tasks, such as image-grounded reasoning.
Is there a repository that links to any or all papers or systems that use the dataset?
Answer: No.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
Answer: The annotations of the MMR benchmark are manually annotated by us, whereas the annotations of MMR-data are generated by GPT-4o. All metadata will be publicly accessible in the dataset repository. The images are collected from websites or open-source datasets. We have taken measures such as changing the resolution, modifying the aspect ratio, changing the format, and editing to minimize the impact on the original work rights. However, it is still possible that the copyright holder may request the deletion of certain data. If this happens, we will edit the content without affecting the questions and answers. If retaining images is also not allowed, we will still keep the annotation data and provide metadata (including URL) for the corresponding images.
Are there tasks for which the dataset should not be used?
Answer: The MMR benchmark and MMR-data cannot be used to evaluate the image or video generation capabilities of the MLLM.
E.5 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Answer: Yes. The MMR benchmark and MMR-data are publicly available on the Internet.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
Answer: The MMR benchmark and MMR-data are available on GitHub https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark or on Huggingface at https://huggingface.co/datasets/BAAI/Multimodal-Robustness-Benchmark.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
Answer: CC BY 4.0.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
Answer: No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
Answer: No.
E.6 Maintenance
Who will be supporting/hosting/maintaining the dataset?
Answer: The authors will be supporting, hosting, and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Answer: Please contact one of the authors (yliu292@connect.hkust-gz.edu.cn; 1ce@bupt.edu.cn; bozhaonanjing@gmail.com)
Is there an erratum?
Answer: No. We will make announcements if there are any.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
Answer: People may appear in the reference images. People may contact us to exclude specific data instances if they appear in the reference images.
Will older versions of the dataset continue to be supported/hosted/maintained?
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
Answer: If others wish to add data, they can apply to do so provided the data is compliant and reasonable. However, making other modifications based on our dataset is currently not allowed.